Grade 10Trigonometry

prove the following cot2A.cot2B+cot2B.cot2C+cot2C.cot2A=1

aryan pokharl
8 Years agoGrade 10

1 Answer

Arun
8 Years ago
Dear Aryan
 
a + b + c = π; so (a + b) = π - c 
==> cot(a + b) = cot(π - c) = -cot(c) 

==> {cot(a)*cot(b) - 1}/{cot(a) + cot(b)} = -cot(c) 
Simplifying the above, 
cot(a)*cot(b) + cot(b)*cot(c) + cot(c)*cot(a) = 1 
 
 
Regards
Arun(askIITians forum expert)